the volume of a right circular cylinder is 96m cubed.express this volume in cubic millimetres, using scientific notation.
Answer by greenestamps(13200)



1 m^3 = (1000mm)^3 = (10^3 mm)^3 = 10^9 mm^3

96 m^3 = 96*10^9 mm^3.

In correct scientific notation,
96*10^9 mm^3 = (9.6*10)*10^9 mm^3 = 9.6*10^10 mm^3.

Answer: 9.6*10^10 mm^3.